# Speculative Knowledge Distillation: Bridging the Teacher-Student Gap Through Interleaved Sampling

### v1

@@ Appendix E Task input and output Format
Appendix F Evaluation Metrics
@@
Tables 7 and 8 illustrate the structured output format used to derive final answers from math instruction tasks and GSM-8k. We employed the evaluation code from (Yuan et al., 2024) to assess the mathematical equivalence of parsed answers and ground truths. This involved converting mathematical symbols like $\pi$ into their numerical values for more accurate numerical comparisons, rather than relying solely on string matching. Implementation details can be found in the original paper (Yuan et al., 2024). We reported accuracy for both tasks.
@@
Appendix G Case Study between On-policy and SKD samples
Tables 11 and 12 illustrate low-quality training examples generated by on-policy models. In Table 11, the student model hallucinates ”Estonian language,” a nonexistent term in the source, leading to a poor translation. SKD with teacher’s interleaved sampling avoids this issue. In Table 12, on-policy KD samples repeat the n-gram ”in the past,” producing nonsensical translations. SKD’s interleaved sampling helps prevent this pathology. In Table 13, we sampled training data from SKD and on-policy methods using the Gemma-2B-it model. These samples were collected during the early stages of training. The on-policy sample exhibited repetitive outputs and errors in arithmetic calculations. While our SKD model also produced incorrect answers, incorporating teacher’s corrections enabled it to perform some intermediate steps correctly and maintain accurate arithmetic calculations. In Table 14, we sampled training data from SKD and on-policy methods using the Gemma-2B-it model. These samples were collected during the late stages of training. We observed that teacher model can perform some intermediate correction on the reasoning step, ultimately leading to correct answer.
@@ Appendix G Case Study between On-policy and SKD samples
Appendix H Over-fitting at SFT stage can cause post-training KDs to sub-optimal performance
@@
Appendix I Training details of mixed training of supervised KD and on-policy KD
The parameter setting for this baseline is identical for the setup in Appendix C. We start to train student model with supervised KD samples for $188$ steps (roughly $1.5$ epoches), then we perform on-policy KD for another $187$ steps. Total training step is $375$. Initial learning rate is $1e$-$5$ and linear decay is applied during training process.
Appendix J speedup to speculative decoding

### v2

@@ Appendix F Task input and output Format
Appendix G Evaluation Metrics
@@
Tables 8 and 9 illustrate the structured output format used to derive final answers from math instruction tasks and GSM-8k. We employed the evaluation code from (Yuan et al., 2024) to assess the mathematical equivalence of parsed answers and ground truths. This involved converting mathematical symbols like $\pi$ into their numerical values for more accurate numerical comparisons, rather than relying solely on string matching. Implementation details can be found in the original paper (Yuan et al., 2024). We reported accuracy for both tasks.
@@
Appendix H Case Study between On-policy and SKD samples
Tables 12 and 13 illustrate low-quality training examples generated by on-policy models. In Table 12, the student model hallucinates ”Estonian language,” a nonexistent term in the source, leading to a poor translation. SKD with teacher’s interleaved sampling avoids this issue. In Table 13, on-policy KD samples repeat the n-gram ”in the past,” producing nonsensical translations. SKD’s interleaved sampling helps prevent this pathology. In Table 14, we sampled training data from SKD and on-policy methods using the Gemma-2B-it model. These samples were collected during the early stages of training. The on-policy sample exhibited repetitive outputs and errors in arithmetic calculations. While our SKD model also produced incorrect answers, incorporating teacher’s corrections enabled it to perform some intermediate steps correctly and maintain accurate arithmetic calculations. In Table 15, we sampled training data from SKD and on-policy methods using the Gemma-2B-it model. These samples were collected during the late stages of training. We observed that teacher model can perform some intermediate correction on the reasoning step, ultimately leading to correct answer.
@@ Appendix H Case Study between On-policy and SKD samples
Appendix I Over-fitting at SFT stage can cause sub-optimal performance at post-training KDs
@@
Appendix J Training details of mixed training of supervised KD and on-policy KD
The parameter setting for this baseline is identical for the setup in Appendix D. We start to train student model with supervised KD samples for $188$ steps (roughly $1.5$ epoches), then we perform on-policy KD for another $187$ steps. Total training step is $375$. Initial learning rate is $1e$-$5$ and linear decay is applied during training process.
Appendix K speedup to speculative decoding
@@
Appendix L Rejection Rate Analysis
We calculate the rejection rate of the teacher model across different student initializations and training steps. The rejection rate is defined as the total number of tokens re-sampled by the teacher divided by the total number of tokens generated by both the teacher and student. After 1,225 training steps, the rejection rate decreases to 5.4% for the instruction-tuned Gemma-2B and 0.1% for the supervised fine-tuned (SFT) Gemma-2B, respectively.
We randomly select 100 samples from the Assamese-to-English training dataset. Figure 9 illustrates the estimated rejection rates of the teacher model (Supervised Fine-Tuned Gemma-7B) with two student models: Gemma-2B-IT and Gemma-2B-SFT. For Gemma-2B-IT (left panel), the rejection rate starts high, with 40.9% of tokens requiring teacher resampling during sequence generation. This rate drops sharply within the first 150 training steps, from 40.9% to 9.7%, and continues to decline gradually as training progresses. This trend supports our hypothesis that, in early training stages, SKD behaves more like supervised knowledge distillation (with  40.9% of tokens replaced by the teacher). As training advances, SKD transitions to a more on-policy approach, with only  5% of tokens replaced by the teacher.
In the right panel of Figure 9, we evaluate the rejection rate with the SFTed Gemma-2B student model. The initial rejection rate is significantly lower (4.7%), reflecting the student model’s stronger initialization and closer proximity to the teacher. The rejection rate continues to decrease as training progresses, maintaining a consistent downward trend. Despite the lower initial rejection rate and the more on-policy behavior of SKD with the SFTed student model, we demonstrate SKD’s superior performance in Table 2.
Figure 9: We compute the rejection rate of the teacher model under different student initializations with various training steps at Assamese to English translation task. We estimated rejection rates of teacher model supervised FT Gemma-7B with student models gemma-2b-it and gemma-2b-sft, respectively. We found that the rejection rate has a sharp decrease in the first $150$ iterations. Student model with higher quality initialization (SFTed) can achieve much lower rejection rate.
Appendix M Computational Analysis
An important consideration in SKD is the computational cost associated with on-policy sampling. Using the estimated rejection rates as a reference, we can analyze the computational costs of this approach. By leveraging the isoFLOP cost calculation introduced in (Kaplan et al., 2020), we estimate the inference cost of our models using Equation 2:
$Computation=2ND+2N\sqrt{D}*L$ (2)
where $N$ is the number of parameters. $D$ is the number of input tokens and $L$ is the number of decoding steps.
Based on our rejection rate analysis, we compute the expected rejection rate for Gemma-2B-IT during training. For simplicity, we assume an expected rejection rate of 10% (slightly higher than the actual rate during training but convenient for analysis), corresponding to an expected acceptance rate of 90%. We estimate the expected number of generated tokens using a simplified version of Algorithm 1 shows a simplified algorithm where $\gamma=1$. In practice, we use $\gamma=5$.
Using a geometric series approach similar to (Leviathan et al., 2022), we estimate the expected number of student-generated tokens with Equation 3:
$E(\#\ student\ generated\ tokens)=\frac{1-\alpha^{\gamma+1}}{1-\alpha}$ (3)
where $\alpha$ is the expected acceptance ratio and $\gamma$ is the sequence length generated by the student model per run. With an expected acceptance ratio of $0.9$, the expected number of student-generated tokens per run is $(1-0.9^{5})/(1-0.9)=4.1$. For simplicity, we round this to $4$. For $5$ student-proposed tokens, the teacher evaluates all tokens in parallel and typically rejects the last one. Thus, we can approximate that in SKD, the student generates 80% of the tokens, while the teacher generates the remaining 20%.
To quantify this in terms of inference cost, we assume $N_{student}$=$2B$, $N_{teacher}$=$7B$, $D$=$256$, $L$=$200$. If the ground truth is unavailable and must be sampled from the teacher, supervised KD is roughly 3.5x higher than on-policy KD, while the sampling cost of SKD is approximately 1.72x higher than on-policy KD. SKD can thus save approximately 50% of the computational cost compared to directly sampling from the teacher. This calculation is based on an instruction-tuned student model; using an SFTed student model would yield additional computational savings. This example provides a high-level overview of the computational requirements for SKD sampling.
Appendix N Validation Convergence across methods
Figure 10: In this Figure, we report the validation loss of various baseline KD strategies across 350 training steps, distilling Gemma-7B to Gemma-2B-it model. Our findings demonstrate that SKD consistently achieves the lowest validation loss by the final iteration across all three tasks.
Figure 11: We report the validation loss of various baseline KD strategies across 350 training steps, distilling Gemma-7B to SFTed Gemma-2B model. Our findings demonstrate that SKD consistently achieves the lowest validation loss by the final iteration across all three tasks.
In Figure 10, we present the validation loss of various baseline KD strategies over 350 training steps, where Gemma-7B is distilled into the Gemma-2B-IT model. Our results show that SKD consistently achieves the lowest validation loss by the final iteration across all three tasks. A key observation is that the lower quality of on-policy samples can lead to slower initial convergence. This is evident at step 50, where on-policy KD exhibits higher validation losses compared to both supervised KD and SKD across all tasks. In translation tasks, student models trained with on-policy KD continue to show higher validation losses than those trained with SKD or supervised KD, reflecting the inferior quality of on-policy samples, as indicated by the lower COMET scores in Table 1. For simpler tasks, such as translation and summarization, supervised KD achieves lower initial validation losses compared to SKD and on-policy KD. However, SKD rapidly improves in later iterations, ultimately achieving the lowest validation loss by the end of training, while supervised KD risks overfitting, as shown in the rightmost panel of Figure 10. For more complex tasks, such as the GSM task, SKD’s interleaved sampling method enables faster convergence and consistently lower validation losses throughout training, outperforming both on-policy and supervised KD.
In Figure 11, we report the validation loss of baseline KD strategies over 350 training steps, distilling Gemma-7B into the SFTed Gemma-2B model. SKD consistently achieves the lowest validation loss by the final iteration across all three tasks. Additionally, we observe that as the base quality of the student model improves, on-policy KD achieves lower validation losses compared to supervised KD by the end of training. This suggests that on-policy approaches benefit from improved sample quality. Most importantly, SKD achieves the best convergence within 350 training steps for both model initializations, outperforming both supervised KD and on-policy KD.
The training process remains stable across on-policy KD, supervised KD, and SKD, as the KL loss provides a stable training objective. The primary difference between these approaches lies in the effectiveness of the samples for learning, as measured by their ability to minimize validation loss. Through a comparative analysis of validation losses for both instruction-tuned (IT) and supervised fine-tuned (SFT) student model initializations, we demonstrate that SKD achieves superior convergence, consistently attaining the lowest validation loss compared to baseline methods.
Appendix O Quantify sample OOD issues with different KD approaches
To quantify the out-of-distribution (OOD) issues in the samples, we computed the per-sample perplexity (PPL) using both the teacher model (supervised fine-tuned Gemma-7B) and the student model (instruction-tuned Gemma-2B). Our analysis included ground truth data, on-policy samples generated by the student, and samples produced by SKD. We randomly selected 100 prompts from the translation training dataset for student and SKD sampling, while for supervised KD, we used ground truth outputs.
As shown in Table 16, the teacher model assigned a significantly higher PPL to student-proposed samples $46.8$, compared to ground truth samples $1.57$, indicating a substantial divergence between the student’s proposals and the teacher’s distribution. In contrast, SKD-generated samples achieved a much lower PPL of $10.8$, suggesting that the teacher model can more accurately evaluate the quality of SKD samples compared to those proposed by the student.
Table 16: In this Table, we quantify sample out-of-distribution (OOD) issues by estimating the per-sample perplexity (PPL), using both the teacher (supervised FT Gemma-7B) and the student (instruction-tuned Gemma-2B). Our analysis included ground truth data (supervised KD), on-policy samples proposed by the student, and samples generated by SKD’s interleaved sampling. Our finding demonstrates that both student proposed samples and ground truth are OOD to teacher and student model respectively. With large drops at PPL value, we demonstrate that SKD’s interleaved sampling can mitigate sample OOD issues and lead to more effective KD process.
PPL from SFTed Gemma-7B (Teacher) PPL from Gemma-2B-it (Student)
On-policy SKD Supervised KD On-policy SKD Supervised KD
PPL 46.8 10.8 1.57 44.0 358 5606.3
Table 16 further reveals that ground truth samples exhibit a high PPL of 5606.3 under the student model, confirming their off-policy nature. Importantly, our SKD approach, which combines on-policy token proposals with teacher corrections, significantly reduces the PPL to 358 compared to ground truth samples. This demonstrates that SKD samples are better aligned with the student model’s distribution, enabling more effective on-policy learning.